OpenStudy (anonymous):

If m - n = 32 and 3m - 8n - 6 = 0, then 4m - 9n = a. -8 b.18 c.26 d.38 e.50

OpenStudy (anonymous):

so here it is.... solve one of the equations for m and the other for n. Then substitute it into the other equation. so say we solve equation 1 for m = 32 +n solve equation 2 for n = (6-3m)/8 now substitute and solve.

OpenStudy (anonymous):

now multiply m-n=32 with 3 and solve it with 3m-8n=6 then you will get n=18 sub it in m-n=32 now m=50 now 4m-9n= 4(50)-9(18)=200-162=38